OpenStudy (anonymous):

Sketch the region given by the set {(x,y)|4 <= x <= 5,-5<=y<=1} on a piece of paper. The area of the region is [blank] square units. Help please!

OpenStudy (anonymous):

see attached graph

OpenStudy (anonymous):

How would I put the answer numerically? Like it just has one blank and it says what's the area of the region in square units? As I have shown it in the question up top.

OpenStudy (anonymous):

the width in the x-direction is 5-4=1 the length in the y-direction is 1-(-5)=6 so the area=1*6=6 square units

OpenStudy (anonymous):

The system took it as a correct answer! Thanks!

OpenStudy (anonymous):

np :})